A library expansion, began in 2002, was expected to cost $103 million. by 2006, library officials estimate the cost would be $29 million over budget. By what percent did the cost exceed the initial estimate?
Answer by checkley79(3341)

(103+29)/103=132/103=1.28 OR 28% OVER ESTIMATE.